OpenStudy (anonymous):

Eric can paint a wall in 20 minutes. His brother, Jessup, can paint the same wall in 30 minutes. How long will it take them working together to paint the wall?

OpenStudy (anonymous):

y=2/3x where y = the time it takes eric

OpenStudy (anonymous):

\[\frac{20\times 30}{20+30}\]

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Do you speak spanish??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Eric 1 wall - 20 minutes X xall - 1 minute

OpenStudy (anonymous):

then X = 1/20 wall

OpenStudy (anonymous):

the same forJessus

OpenStudy (anonymous):

Jessus : Y=1/30 wall

OpenStudy (anonymous):

in one minute so if they work together, you just have to....(i don't know the word) add? X+Y

OpenStudy (anonymous):

and that´s waht they both do in one minute

OpenStudy (anonymous):

why is a good question. lets work it out. the entire job is one job, eric does \[\frac{1}{20}\] per minute and his friend does \[\frac{1}{30}\] per minute so their combined rate is \[\frac{20+30}{20\times 30}\] and you want to solve \[\frac{20+30}{20\times 30}T = 1\] and therefore you get \[T=\frac{20\times 30}{20+30}\] and now you never have to do it again

OpenStudy (anonymous):

i need help with my homework.

OpenStudy (anonymous):

erick does the job in x hours, joe does it in y hours their combined rate is \[\frac{1}{x}+\frac{1}{y}=\frac{x+y}{xy}\] and therefore the job gets finished by solving \[\frac{x+y}{xy}T = 1 \] thus \[T=\frac{xy}{x+y}\]

OpenStudy (anonymous):

thank yall.